OpenStudy (anonymous):

Help me!! Will fan and give medal Let f(x) = 5x + 12. Find f-1(x). 12x + 5 1 over quantity 5 x plus 12 5 over quantity x minus 12 quantity of x minus 12 over 5

OpenStudy (anonymous):

@jose12345

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

By f-1(x) you mean the inverse I assume. To do that just swap the values of your function with y<->x. y=5x+12 becomes x=5y+12 and now solve for y. Do you know how to solve the equation for y, that means to rearrange the equation I have given, x=5y+12 to have y alone and all the numbers on x's side. you need to do algebraic operations. To get 12 on the other side, you need to subtract 12 from both sides, so you will have, x-12 =5y Now what do you do with the 5 that is multiplying y. How do you move that from the right side to the left side? I just wanted to know how to explain the problem to you. Alright, check this out. x-12=5y You divide both sides by 5 because that number is multiplying y on the right side. So when you divide both sides you will cancel out the one that's multiplying y and. Therefore your equation now becomes, x/5 - 12/5 = y That is the inverse of your initial function if you had something like 1/5x - 12/5, that would be the closest

OpenStudy (anonymous):

So whats the answer

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

smart

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

who me

OpenStudy (anonymous):

Answer?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

thats for me can u answer my message

OpenStudy (anonymous):

cool i will

OpenStudy (anonymous):

brb im going to do the question

OpenStudy (anonymous):

is b